# A Comprehensive Framework to Operationalize Social Stereotypes for Responsible AI Evaluations

### v2

@@ 4 Framework
The cognitive process of categorizing encourages people to think in terms of “us” (in-group) vs. “them,” (out-group) which in turn leads to stereotyping. The out-group, or target group, is an integral component of a stereotype, and can be characterized with the following features:
@@
  • Intersectional. Theories of social categorization explain that perceiving an individual as a member of multiple groups (either considered as the perceiver’s in-groups or out-groups) leads to specific stereotypes beyond the ones associated to either of the constituent groups. The perceiver’s judgement might change when they categorize the target into in-group gender but out-group race, as opposed to out-group gender and race. So whether the group is intersectional or not is an important aspect to capture.
  • Marginalized. If a social group is historically marginalized, stereotypes portent to result in more harm. This may result in discriminatory hiring practices enabled by AI systems magnifying stereotypes about temperament and suitability of employment about women and African-Americans who are known to have been marginalized in the US Bertrand and Mullainathan (2004); Chen (2023). Capturing such historical marginalization may help determine (and prioritize) the appropriate course of action once stereotypes are detected in model output.
  • Demographic. A social group can be defined by demographic features such as race, gender, or age, and other extrinsic or acquired attributes such as profession or life style. Non-demographic groups may be more fluid and self-selected, whereas demographic groups are based on fixed or inherent characteristics. Stereotypes about demographic groups are often intertwined with social dynamics and can be associated with systematic discrimination. Therefore, it is important to capture this distinction.
@@
The attribute describes the beliefs, assumptions, features, sentiments, or perceptions that are widely associated with members of the target group. While the association of these attributes to the target group is core to the notion of stereotypes, the attributes themselves can be characterized with certain features:
  • Valence. The valence of the attribute can include aspects such as the associated perceived offensiveness Jha et al. (2023), warmth, competence Nicolas et al. (2021), or morality Fiske et al. (2018) of the term. The perceptions of attributes as such, and what motivates people to use them, is discussed in social psychology and NLP literature and can inform practices that rely on human ratings for identifying stereotypes. The valence of attributes may also help NLP practitioners prioritize debiasing efforts (e.g., focusing on stereotypes with offensive attributes).
  • Modality. Attributes manifest in different ways across different modalities. For instance, attributes like ‘soft spoken’ or ‘intelligent’ can be expressed clearly in text, video, or audio, but less likely to be depicted in images. On the other hand, the markers of ‘poverty’ can be vastly different in text (e.g., descriptions of poverty) versus image or video (e.g., dusty streets as visual markers of poverty that are not often verbalized). Capturing this nuance is crucial to operationalize such large databases of socio-cultural information into robust model or data interventions.
@@
  • Statistical Basis (cf. Accuracy). The distinction between whether an association is a stereotype or factual/definitional is often blurry. For example, while it is true that Hindus often pray in temples, and this association is statistically accurate, generalizing all Hindus as temple-goers can be perceived as stereotyping, as Hinduism (like any religion) in practice encompasses a wide range of rituals beyond temple worship. On the other hand, certain associations may be readily accepted as stereotypes, but also has statistical basis: for instance, some occupational stereotypes found in NLP models align with actual US census data on job distribution Garg et al. (2018b).
@@
  • Salience or Prevalence. The salience or prevalence of the association can be described in various levels. It is useful to distinguish them at least at two levels from an NLP perspective: (1) model/data/language salience represents how frequently or prominently the association appears in the model or dataset in a given language and can be measured in different ways Jha et al. (2023); Bhatt et al. (2022). Model salience can further be an indicator of how likely it is to influence model generations. (2) Social salience captures how widespread an association is in society, captured either at a global level, or variations across regions and communities.
@@
The stereotypical association of the attribute with the target group is held by a group of people or a section of society, who are the perceivers of the stereotype. The socio-economic standing of this group of people, and the fraction of population they account for are some aspects that may be of relevance.
  • Social Group. The social group that the perceivers belong to is crucial in understanding stereotypes because it significantly influences how they distinguish in-groups from out-groups and consequently perceive and interact with the target group. It is also important to note that any implications of social group membership of perceivers will differ from those of target group’s social axes. For instance, whether or not a target group is historically marginalized may be crucial in determining how stereotypes about them may be prioritized in certain contexts, but whether the perceiver group was historically marginalized or not may not be hold the same weightage.
  • Region/Social context. Social groups often have different levels of power and status in society. This power differential can also influence how stereotypes are formed and perpetuated. Therefore the interaction of perceivers’ social group and the target group is meaningful in this context. This dynamic is an important factor for determining the possible harmfulness of the stereotype.
@@
Finally, it is crucial to remember that stereotypes are not universal or static. They exist within specific social, cultural, and temporal contexts. Instead of implying that we are speaking about “society” in general, it is important to pinpoint both the time period and the specific reference/artifact (a dataset, a model, a geo-cultural region, etc.) that reflects the societal views in question. This precision will help avoid generalizations and ensures a more accurate analysis of stereotypes.
@@
  • Reference. Stereotypes captured in NLP datasets and models, exist within specific socio-cultural contexts. Their prevalence may vary depending on which slice of society is captured in any specific dataset or model. Hence, it is important to also capture this referential context – i.e., which societal context, and which artifact, whether data or model.
@@
It is important to also note that the features in the framework may interact with one another. For example, Christians are a minority group in India and can be seen as marginalized, whereas, the same group is not similarly marginalized in the US. This difference may influence how stereotypes about the same target group may be dealt with in India vs. the US.
@@ 5 Roadmap for Operationalization
The framework presented above is intentionally broad, with the aim to capture all aspects of stereotypes that may be relevant in responsible AI evaluations. There may be crucial considerations that helps when it comes to operationalizing the framework in specific contexts. In this section we provide such a roadmap for implementation and utilizing the framework.
@@ 5.1 Recommendations for Implementation
[…]
While such a simplistic implementation may suffice for demonstrative purposes, and for small scale evaluations, most real-world scenarios will require a more sophisticated implementation that can account for inter-relationships between various elements of the framework. In particular, a knowledge graph based implementation might be especially appropriate in this case, as it will support sophisticated analytics for robust data exploration and visualization, a high level of expressiveness to capture complex contextual and metadata details, adaptability to accommodate evolving insights about stereotypes, and extensibility to incorporate related entities and information from other resources.
Knowledge graphs allow for flexible data modeling Angles et al. (2017), which is crucial for capturing the evolving nature of stereotypes and their associated attributes. They emphasize relationships, enabling modeling complex relationships Paulheim (2017) between stereotypes and other components such as social groups. Knowledge graphs also enable capturing nuanced knowledge about context, such as time, locale, and source provenance associated with stereotypes. Their semantic capabilities enable automated reasoning and insights, with structures suited to complex queries, visualization, and pattern detection Hogan et al. (2021). Stored in graph databases, knowledge graphs support rapid data retrieval and efficient scaling, aided by query optimization techniques like partitioning and indexing Angles et al. (2017), making them ideal for downstream mitigation efforts.
@@ 5.2 Utilizing the Framework
In this section, we outline some of the ways in which our framework bridges many of of the gaps identified in Section 3.3. Which of those gaps are crucial may vary on a case by case basis. For instance, if an evaluation is aimed to be applied in a mono-lingual mono-cultural setting, then the geo-cultural specification on stereotypes’ context may not be crucial in that case.
@@
  • Targeted evaluation. Our framework can facilitate verifying whether model’s responses contain specific categories of stereotypes. For instance, one might be interested in stereotypes involving identities related to a particular social axis, such as race, religion, or nationality where the identity might be that of the target group or the perceiver; stereotypes where the target is a marginalized group; stereotypes that are particularly offensive in some context; stereotypes that are prevalent in a particular culture and/or region; and more. A unified framework lends itself for such comprehensive and targeted evaluations.
@@
  • Temporal evolution analysis. The temporal dimension in our framework allows us to track how the prevalence, valence, and/or social groups associated with stereotypes have changed over time. For instance, it was shown that gender stereotypes have evolved over time Garg et al. (2018b), with newer stereotypes emerging in different periods of time. Similarly, on evaluation of stereotypes and associated offensiveness, general trends of perception of different groups of people can be determined.
@@
  • Differences. We can analyze stereotypes associated with a particular group according to different perceivers. This might be useful to understand how groups along a given spectrum may perceive a certain relevant group to gauge deeper concerns that perceivers might have about the target. For instance, we could compare the stereotypes held by democrats and republicans in the US towards certain groups of people, such as immigrants, trans people, or atheists.
@@
Preventing Silowed Evaluations with Stereotype Interdependency:
@@
  • Co-occurrence analysis. Stereotypes can frequently co-occur, and magnify different aspects of marginalization, such as stereotypes about race and gender, or social class and ethnicity Bond et al. (2021). Such patterns reveal important inter-dependencies that our framework enables us to identify in data and models, which in-turn could lead to preventing harms to intersectional groups.
  • Conflict and Synergy analysis. Multiple stereotypes can exist in a society such that they conflict or contradict each other, leading to social tensions (e.g., immigrants as both lazy and stealing jobs). Stereotypes may also coexist and thus can reinforce or amplify one another, creating a more harmful impact, for instance, black women being stereotypes as loud and angry, can lead to workplace discrimination Motro et al. (2021). This framework enables analysis and aggregation of such interdependencies at local and global scales.
@@
Understanding how stereotypes emerge and spread is essential for developing effective interventions, and our framework provides the tools for tracing these patterns.
@@
  • Media analysis. The media often plays a critical role in shaping the perception of people worldwide, [^2] https://www.chicano.ucla.edu/files/news/NHMCLatinoDecisionsReport.pdf and in turn it also captures and reinforces perceptions of people already present in society. [^3] https://blog.google/intl/en-in/company-news/using-ai-to-study-demographic-representation-in-indian-tv/ Analyzing media representations, such as movies, television shows, or news articles, contribute to understanding of the formation and/or reinforcement of stereotypes.
@@ 6 Discussion
Stereotypes are known to be prevalent in everyday societal interactions, which in turn find their way into the massive datasets used to train language models, and are often reflected in those model predictions. While researchers have made efforts to evaluate and mitigate this issue, existing approaches on this topic often lack a cohesive structure and tend to focus on isolated aspects of the problem. In this paper, we introduced a comprehensive framework for systematically evaluating stereotypes in language technologies, focusing on five core aspects: target group, attribute, association, perceiver, and context, each with an associated set of features. We also outline detailed guidelines for implementing and utilizing this framework in practice. By offering a holistic and operationalizable approach, we aim to empower researchers and developers to develop responsible AI approaches towards language technologies that are effective and robust.

### v3

@@ 4 Framework
The cognitive process of categorizing encourages people to think in terms of “us (in-group) vs. them (out-group),” which in turn leads to stereotyping. The out-group, or target group, is fundamental to stereotype research Allport et al. (1954) and an integral component of a stereotype which can be characterized with the following features:
@@
  • Intersectional. Theories of social categorization explain that perceiving an individual as a member of multiple groups (either considered as the perceiver’s in-groups or out-groups) leads to specific stereotypes beyond the ones associated with either of the constituent groups. The perceiver’s judgment might change when they categorize the target into in-group gender but out-group race, as opposed to out-group gender and race. So whether the group is intersectional or not is an important aspect to capture.
  • Marginalized. If a social group is historically marginalized, stereotypes are more likely to result in more harm. This is not to say that stereotypes toward non-marginalized groups are harmless, rather, the mechanism of harm varies based on whether or not the group is historically marginalized. This may result in discriminatory hiring practices enabled by AI systems magnifying stereotypes about temperament and suitability for employment about women and African-Americans who are known to have been marginalized in the US Bertrand and Mullainathan (2004); Chen (2023). Capturing such historical marginalization may help determine (and prioritize) the appropriate course of action once stereotypes are detected in model output.
  • Demographic. A social group can be defined by demographic features such as race, gender, or age, or other extrinsic or acquired attributes such as profession or lifestyle. Non-demographic groups may be more fluid and self-selected, whereas demographic groups are based on fixed or inherent characteristics. Stereotypes about demographic groups are often intertwined with social dynamics and can be associated with systematic discrimination. Therefore, it is important to capture this distinction Crenshaw (2013).
@@
The attribute describes the beliefs, assumptions, features, sentiments, or perceptions that are widely associated with members of the target group. Our conceptualization of the attribute as the characteristic associated with the target group draws heavily from the SCM Fiske et al. (2018); Cuddy et al. (2007). While the association of these attributes to the target group is core to the notion of stereotypes, the attributes themselves can be characterized with certain features:
  • Valence. We directly borrow valence from the SCM; the valence of the attribute can include aspects such as the associated perceived offensiveness Jha et al. (2023), warmth, competence Nicolas et al. (2021), or morality Fiske et al. (2018) of the term. The perceptions of attributes as such, and what motivates people to use them, is discussed in social psychology and NLP literature and can inform practices that rely on human ratings for identifying stereotypes. The valence of attributes may also help NLP practitioners prioritize debiasing efforts (e.g., focusing on stereotypes with offensive attributes).
  • Modality. Attributes manifest in different ways across different modalities. For instance, attributes like “soft spoken” or “intelligent” can be expressed clearly in text, video, or audio, but less likely to be depicted in images. On the other hand, the markers of “poverty” can be vastly different in text (e.g., descriptions of poverty) versus image or video (e.g., dusty streets as visual markers of poverty that are not often verbalized). Capturing this nuance is crucial to operationalize such large databases of socio-cultural information into robust model or data interventions.
@@
  • Statistical Basis (cf. Accuracy). The distinction between whether an association is a stereotype or factual/definitional is often blurry. For example, while it is true that Hindus often pray in temples, and this association is statistically accurate, generalizing all Hindus as temple-goers can be perceived as stereotyping, as Hinduism (like any religion) in practice encompasses a wide range of rituals beyond temple worship. On the other hand, certain associations may be readily accepted as stereotypes, but also have statistical basis: for instance, some occupational stereotypes found in NLP models align with actual US census data on job distribution Garg et al. (2018b).
@@
  • Salience or Prevalence. The salience or prevalence of the association can be described in various levels. It is useful to distinguish them at least at two levels from an NLP perspective: (1) model/data/language salience represents how frequently or prominently the association appears in the model or dataset in a given language and can be measured in different ways Bhatt et al. (2022); Jha et al. (2023). Model salience can further be an indicator of how likely it is to influence model generations. (2) social salience captures how widespread an association is in society, captured either at a global level, or variations across regions and communities.
@@
The stereotype is held by a group of people or a section of society, who we refer to as perceivers Turner et al. (1979). By including perceivers into this framework, we acknowledge that stereotypes are not simply properties of target groups but are actively constructed and applied by perceivers—a concept similar to the role of speaker in NLP research Hovy and Yang (2021). The socio-economic standing of this group of people, and the fraction of the population they account for are significant aspects that contribute to the severity of the stereotype.
  • Social Group. The social group that the perceivers belong to is crucial in understanding stereotypes because it significantly influences how they distinguish in-groups from out-groups and consequently perceive and interact with the target group. It is also important to note that any implications of social group membership of perceivers will differ from those of the target group’s social axes. For instance, whether or not a target group is historically marginalized may be crucial in determining how stereotypes about them may be prioritized in certain contexts, but whether the the perceiver group was historically marginalized or not may not hold the same weight.
  • Region/Social context. Social groups often have different levels of power and status in society. This power differential can also influence how stereotypes are formed and perpetuated. Therefore the interaction of the perceivers’ social group and the target group is meaningful in this context. This dynamic is an important factor for determining the possible harmfulness of the stereotype.
@@
Finally, it is crucial to remember that stereotypes are not universal or static. They exist within specific social, cultural, and temporal contexts that shape human behaviors Lewin (1951). Instead of implying that stereotypes speak about “society” in general, it is important to pinpoint both the time period and the specific reference/artifact (a dataset, a model, a geo-cultural region, etc.) that reflects the societal views in question. For instance, the perceived social norms and support for prejudice reduction in a given context can influence whether people express prejudiced attitudes Devine and Elliot (1995). This precise component will help prevent generalizations and ensures a more accurate analysis of stereotypes.
@@
  • Reference. Stereotypes captured in NLP datasets and models, exist within specific socio-cultural contexts. Their prevalence may vary depending on which slice of society is captured in any specific dataset or model. Hence, it is important to also capture this referential context—i.e., which societal context, and which artifact, whether data or model.
@@
It is important to also note that the features in the framework may interact with one another. For example, Christians are a minority group in India and can be seen as marginalized, whereas, the same group is not similarly marginalized in the US. This difference influences how stereotypes about the same target group may be dealt with in India vs. the US Kulkarni et al. (2023).
@@ 5 Roadmap for Operationalization
The framework presented above is intentionally broad, with the aim of capturing all aspects of stereotypes that may be relevant in responsible AI evaluations. There may be crucial considerations that help when it comes to operationalizing the framework in specific contexts. In this section we provide such a roadmap for implementation and utilizing the framework.
@@ 5.1 Recommendations for Implementation
Our framework is conceptual in nature, and is not tied to any particular implementation approach. A simpler implementation, for instance, using spreadsheets or relational databases, may suffice if the evaluation context is narrowly scoped. Table 1 shows one such tabular form implementation of our framework, where we mapped instances from five stereotype resources that are prominently used in NLP. We chose approximately 20 examples from each of the datasets, and mapped the existing information in those datasets onto our framework. This exercise revealed cases where certain features are not applicable (e.g., vegetarianism as an attribute does not lend itself to the SCM categories of Warmth and Competence, as it is based on a religious practice. It also revealed cases where existing datasets lack certain relevant information; e.g., StereoLMs dataset does not capture perceiver information, whereas SeeGULL and SPICE capture regional information of perceivers.
While such a simplistic implementation may suffice for demonstrative purposes, and for small scale evaluations, most real-world scenarios will require a more sophisticated implementation that can account for interrelationships between various elements of the framework. In particular, a Knowledge Graph-based implementation might be especially appropriate in this case, as it will support sophisticated analytics for robust data exploration and visualization, a high level of expressiveness to capture complex contextual and metadata details, adaptability to accommodate evolving insights about stereotypes, and extensibility to incorporate related entities and information from other resources.
Knowledge Graphs allow for flexible data modeling Angles et al. (2017), which is crucial for capturing the evolving nature of stereotypes and their associated attributes Deshpande et al. (2022). They emphasize relationships, enabling modeling complex relationships Paulheim (2017) between stereotypes and other components such as social groups. Knowledge Graphs also enable capturing nuanced knowledge about context, such as time, locale, and source provenance associated with stereotypes. Their semantic capabilities enable automated reasoning and insights, with structures suited to complex queries, visualization, and pattern detection Hogan et al. (2021). Knowledge Graphs support rapid data retrieval and efficient scaling, aided by query optimization techniques like partitioning and indexing Angles et al. (2017), making them ideal for downstream mitigation efforts.
@@ 5.2 Utilizing the Framework
In this section, we outline some of the ways in which our framework bridges many of of the gaps identified in Section 3.3. Depending on the use case, researchers should be able to identify which of the mentioned gaps might impact their conceptualization of stereotypes. For instance, if an evaluation is aimed to be applied in a monolingual, monocultural setting, then the geo-cultural specification on stereotypes’ context may not be crucial in that case.
@@
  • Targeted evaluation. Our framework can facilitate verifying whether model responses contain specific categories of stereotypes. For instance, one might be interested in stereotypes involving identities related to a particular social axis, such as race, religion, or nationality where the identity might be that of the target group or the perceiver; stereotypes where the target is a marginalized group; stereotypes that are particularly offensive in some context; stereotypes that are prevalent in a particular culture and/or region; and more. A unified framework lends itself to such comprehensive and targeted evaluations.
@@
  • Temporal evolution analysis. The temporal dimension in our framework allows us to track how the prevalence, valence, and/or social groups associated with stereotypes have changed over time. For instance, it was shown that gender stereotypes have evolved over time Garg et al. (2018b), with newer stereotypes emerging in different periods of time. Similarly, through evaluation of stereotypes and associated offensiveness, general trends of perception of different groups of people can be determined.
@@
  • Differences. We can analyze stereotypes associated with a particular group according to different perceivers. This might be useful to understand how groups along a given spectrum may perceive a certain relevant group to gauge deeper concerns that perceivers might have about the target. For instance, we could compare the stereotypes held by Democrats and Republicans in the US toward certain groups of people, such as immigrants, trans people, or atheists.
@@
  • Generalization. Stereotype tuples are often studied in isolation without their linguistic context. This separation makes it impossible to fully assess the implications of different types of generalizing language Davani et al. (2024). Specifically, effectively identifying harmful language requires understanding the intent behind a generalization, which can range from mere mentioning a bias to actively evoking and promoting it.
Preventing Siloed Evaluations with Stereotype Interdependency:
@@
  • Co-occurrence analysis. Stereotypes can frequently co-occur, and magnify different aspects of marginalization, such as stereotypes about race and gender, or social class and ethnicity Bond et al. (2021). Such patterns reveal important interdependencies that our framework enables us to identify in data and models, which in turn could lead to preventing harms to intersectional groups.
  • Conflict and Synergy analysis. Multiple stereotypes can exist in a society such that they conflict or contradict each other, leading to social tensions (e.g., immigrants as both lazy and stealing jobs). Stereotypes may also coexist and thus can reinforce or amplify one another, creating a more harmful impact, for instance, black women being stereotyped as loud and angry, can lead to workplace discrimination Motro et al. (2021). This framework enables analysis and aggregation of such interdependencies at local and global scales.
@@
Understanding how stereotypes emerge and spread is essential for developing effective interventions Antypas et al. (2024), and our framework provides the tools for tracing these patterns.
@@
  • Media analysis. The media often plays a critical role in shaping the perception of people worldwide, [^2] https://www.chicano.ucla.edu/files/news/NHMCLatinoDecisionsReport.pdf and in turn it also captures and reinforces perceptions of people already present in society. [^3] https://blog.google/intl/en-in/company-news/using-ai-to-study-demographic-representation-in-indian-tv/ Analyzing media representations, such as movies, television shows, or news articles, contributes to the understanding of the formation and/or reinforcement of stereotypes.
Enhancing bias mitigation on NLP models:
  • Bias detection. while common datasets can be used for detecting specific stereotypes in models and text, our framework enables detection on various levels, for example, using our framework, researchers could analyze a large corpus of news articles to detect the prevalence of stereotypes associating marginalized ethnic groups (target group) with offensive words (attribute) within the context of immigration debates (context). This allows for targeted analysis of bias concerning a specific marginalized group within a specific context.
  • Bias mitigation. our framework enables more structured bias mitigation by only focusing on stereotypes with specific tones and levels of harmfulness and impact. Suppose our framework analysis reveals that a language model frequently generates sentences associating black women (intersectional target group) with being emotional (attribute, potentially negative valence and low Competence) in the context of workplace interactions. A bias mitigation strategy could then be designed to specifically target and reduce the frequency of these associations in the model’s output, while perhaps being less concerned with other, less harmful stereotypes.
  • Explainability. The framework can be used to explain the biased behavior of NLP models. For example, if a model makes a biased prediction, the framework can help to identify the underlying stereotypes that might be contributing to the bias.
  • Data Augmentation. The framework can be used to generate counter-stereotypical examples for data augmentation, which can help to improve the robustness and fairness of NLP models. Furthermore, the framework can reveal missing information in datasets, for instance showing that a dataset does not include any information about perceivers or lacks data on intersectional groups.
@@ 6 Discussion
Our framework provides a structured language and ontology that helps the NLP community bridge the gap between social psychological theory and computational operationalization. By forcing the explicit articulation of components like the Perceiver and Context, our model moves stereotype analysis beyond simple (Target, Attribute) tuples. This shift is critical for developing more granular and robust evaluation methodologies that are sensitive to socio-historical nuances. For instance, classifying a bias as merely “racial” is insufficient; a proper evaluation requires specifying the relationship—who is holding the belief (Perceiver) about whom (Target) and in what geo-cultural setting (Context)—to determine the appropriate mitigation strategy. Furthermore, a structure like this is essential for building interdependent stereotype knowledge bases that support complex analytical queries, paving the way for the next generation of context-aware and culture-sensitive debiasing techniques in LLMs.
We provided recommendations for implementing the framework using Knowledge Graphs in Section 5.1, however, we also acknowledge that depending on the specific use case in which stereotypes need to be operationalized, developers might not find it efficient to incorporate all aspects of the framework in their design; for instance, the operational complexity and lack of scalability and the role of human oversight in maintaining such a Knowledge Graph introduce significant costs to a project. Thus Section 5.2 discusses how different research and technical problems benefit from specific aspects of the framework. We also acknowledge the need for research into more computationally lightweight alternatives for implementation that still preserve the framework’s richness, allowing smaller research teams or production systems to adopt its core principles without incurring high maintenance costs.
The current framework provides the what (the components of a stereotype), but future work must integrate the how—specifically, developing methods to parse and encode the linguistic context (e.g., sarcasm, metaphor, active vs. passive voice) that modulates a stereotype’s expression and potential for harm. Future efforts should alos rigorously test and adapt the framework’s components to demonstrate utility in a broader range of NLP tasks beyond LLM evaluation, such as bias detection in knowledge distillation or fairness in multimodal generation. This would solidify the framework’s value as a universal tool for responsible AI.